# Evolutionary Prediction Games

### v1

@@ B.1 Preliminaries
The set of distribution over a set $A$ is denoted by ${\Delta\left({A}\right)}$, and the set of distributions over $[K]$ is denoted by $\Delta^{K}={\Delta\left({[K]}\right)}={\left\{{{\bm{p}}\mid\sum_{k}p_{k}=1,p_{k}\geq 0\ \ \forall k\in[K]}\right\}}$. Following the notations in Sandholm (2010), we denote the space tangent to the $K$-simplex by $T\Delta^{K}={\left\{{z\in\mathbb{R}^{K}\mid\sum_{k}z_{k}=0}\right\}}$. We denote ${\bm{1}}=\left(1,\dots,1\right)\in\mathbb{R}^{K}$, and denote by ${\bm{\Phi}}={\bm{I}}-{\bm{1}}{\bm{1}}^{T}$ the orthogonal projection matrix from $\mathbb{R}^{K}$ to $T\Delta^{K}$.
@@
We denote the space of features by $\mathcal{X}$, the space of labels by $\mathcal{Y}$, the hypotheses class by $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, and the loss function by $\mathcal{L}:\mathcal{Y}^{2}\to\mathbb{R}$. The expected loss of hypothesis $h\in\mathcal{H}$ under population distribution $D\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ is denoted by $\mathcal{L}_{D}(h)=\mathbb{E}_{(x,y)\sim D}{\left[{\mathcal{L}(h(x),y)}\right]}$. A supervised learning algorithm $\mathcal{A}$ is a stochastic mapping from a distribution $D\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ to a classifier $h\sim\mathcal{A}(D)$, formally $\mathcal{A}:{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}\to{\Delta\left({\mathcal{H}}\right)}$.
@@
Each group $k$ is associated with a feature-label distribution $D_{k}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$. For a population state ${\bm{p}}\in\Delta^{K}$, the population mixture distribution is denoted by $D_{\bm{p}}=\sum_{k\in[K]}p_{k}D_{k}$. We use a vector subscript notation to distinguish between group distributions (e.g. $D_{k}$) and mixture distributions (e.g. $D_{\bm{p}}$). For a hypothesis $h\in\mathcal{H}$, the expected loss of group $k$ is denoted by $\mathcal{L}_{k}(h)=\mathbb{E}_{(x,y)\sim D_{k}}{\left[{\mathcal{L}(h(x),y)}\right]}$, and the vector of expected group losses is denoted by $\vec{\mathcal{L}}(h)=\left(\mathcal{L}_{1}(h),\dots,\mathcal{L}_{K}(h)\right)$. For ${\bm{p}}\in\Delta^{K}$, the expected loss of $h$ under distribution $D_{\bm{p}}$ is denoted by $\mathcal{L}_{\bm{p}}(h)=\mathbb{E}_{(x,y)\sim D_{\bm{p}}}{\left[{\mathcal{L}(h(x),y)}\right]}$. The support of a population state ${\bm{p}}$ is denoted by $\operatorname*{support}({\bm{p}})={\left\{{k\in[K]\mid p_{k}>0}\right\}}$.
@@
We follow the notations of Sandholm (2010), and provide the key definitions here for completeness. Population games associate each group $k\in[K]$ with a fitness function $F_{k}({\bm{p}})$, which maps a population state ${\bm{p}}\in\Delta^{K}$ to the evolutionary fitness of group $k$. A Nash Equilibrium (NE) of a population game is a population state ${\bm{p}}^{*}$ which satisfies $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k\in[K]}F_{k}({\bm{p}}^{*})$, and we note that every population game admits at least one Nash equilibrium (Sandholm, 2010, Theorem 2.1.1).
Evolutionary dynamics.
For population dynamics, we make the positive correlation assumption (Sandholm, 2010, Section 5.2), and assume that whenever a population is not at rest, the growth rate of a group is positively correlated with its fitness. Positive correlation implies that all Nash equilibria are rest points of the dynamics (Sandholm, 2010, Proposition 5.2.1). For our simulations, we use the replicator dynamics (Taylor & Jonker, 1978), which satisfy this property. We also note that other dynamics converge towards the replicator dynamics in the limit – For example, Moran in the large-population limit (Traulsen et al., 2006), and multiplicative-weights as the multiplicative factor approaches $1$ and time is renormalized accordingly (e.g., Kleinberg et al., 2011). To simplify analysis, we assume that all groups are present at the onset (${\left|{\operatorname*{support}({\bm{p}}^{0})}\right|}=K$).
Potential games.
@@
A population game is a potential game if it can be represented as a gradient of some function. Formally, $F:\Delta^{K}\to\mathbb{R}^{K}$ is a potential game if there exists $f:\Delta^{K}\to\mathbb{R}$ such that $F=\nabla f$, where $\nabla f$ is the gradient of $f$ over the simplex (see Sandholm, 2010, Section 3.2 for rigorous definitions). We informally state key results that are relevant to our setting: By (Sandholm, 2010, Theorem 3.1.3), a state ${\bm{p}}^{*}$ is a Nash equilibrium of $F$ if and only if it is an extremum of the potential function $f$. By (Sandholm, 2010, Theorem 7.1.2), dynamics satisfying the positive correlation property converge towards Nash equilibria (restricted Nash equilibria in case of replicator dynamics). Finally, by (Sandholm, 2010, Theorem 8.2.1), a population state ${\bm{p}}^{*}$ is a stable equilbrium if and only if it is a local maximizer of $f({\bm{p}})$.
B.2 Basic Claims
B.2.1 Linearity
Proposition B.1 (Expected loss is linear in ${\bm{p}}$ for a fixed hypothesis).
@@ B.2.1 Linearity
B.2.2 Overall accuracy equality
A classifier $h$ satisfies overall accuracy equality if both protected and unprotected groups have equal prediction accuracy. In the context of natural selection, we consider this definition with respect to the groups currently present in the population:
Definition B.2 (Overall accuracy equality; (e.g., Verma & Rubin, 2018)).
Let $h:\mathcal{X}\to\mathcal{Y}$, and let ${\bm{p}}\in\Delta^{K}$ be a population state. $h$ satisfies overall prediction equality if $\mathrm{acc}_{k}(h)=\mathrm{acc}_{k^{\prime}}(h)$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}})$.
Under retraining, a different classifier is deployed at each time step. It is therefore natural to define this property with respect to the learning algorithm, rather than a single classifier. In the definition, we take the expectation over the stochasticity of the sampling and learning process to representing the average over time under retraining:
Definition B.3 (Overall accuracy equality in expectation).
Let ${\bm{p}}$ be a population state, and let $\mathcal{A}({\bm{p}})$ be a learning algorithm. $\mathcal{A}$ satisfies overall prediction equality in expectation if $\mathbb{E}_{h\sim\mathcal{A}({\bm{p}})}{\left[{\mathrm{acc}_{k}(h)}\right]}=\mathbb{E}_{h\sim\mathcal{A}(D)}{\left[{\mathrm{acc}_{k^{\prime}}(h)}\right]}$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}})$.
Proof of Proposition 3.2.
Let ${\bm{p}}^{*}$ be a Nash equilibrium of the evolutionary prediction game induced by $\mathcal{A}({\bm{p}})$. By Equation 3, it holds that $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k}F_{k}({\bm{p}}^{*})$, and therefore there exists some $a\in\mathbb{R}$ such that $F_{k}({\bm{p}}^{*})=a$ for all $k\in\operatorname*{support}({\bm{p}}^{*})$. Then by the definition of the evolutionary prediction game (Definition 3.1), it holds that: $\mathbb{E}_{h\sim\mathcal{A}({\bm{p}})}{\left[{\mathrm{acc}_{k}(h)}\right]}=\mathbb{E}_{h\sim\mathcal{A}(D)}{\left[{\mathrm{acc}_{k^{\prime}}(h)}\right]}=a$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}}^{*})$. ∎
B.3 Training Once
Definition B.4 (Train-once prediction game).
@@ B.3 Training Once
Definition B.5 (Constant game; (Sandholm, 2010, Section 3.2.4)).
@@
Fix $h_{0}\in\mathcal{Y}^{\mathcal{X}}$. In the train-once setting, the expected marginal accuracies are constant as a function of ${\bm{p}}$, and therefore $F_{k}$ is a constant game by Definition B.5. By (Sandholm, 2010, Proposition 3.2.13), $F_{k}$ a potential game with linear potential $f({\bm{p}})=-\vec{\mathcal{L}}(h^{0})\cdot{\bm{p}}$. By (Sandholm, 2010, Theorem 3.1.3), the set of Nash equilibria is the set of local maximizers of $f({\bm{p}})$, which is a convex combination over the set $\operatorname*{arg\,min}_{k\in[K]}\mathcal{L}_{k}(h^{0})$ due to linearity. Hence:
@@
At equilibrium ${\bm{p}}^{*}$, it holds that $\mathcal{L}_{{\bm{p}}^{*}}=\min_{k\in[K]}\mathcal{L}_{k}(h^{0})$. Apply Proposition B.1 to obtain:
@@
Finally, observe that each ${\bm{p}}^{*}$ is stable as a maximizer of the linear potential function (Sandholm, 2010, Theorem 8.2.1). ∎
Proof of Proposition 4.1.
Proposition 4.1 is a special case of Proposition B.6 for the loss function: $\mathcal{L}_{k}(h)=-\mathrm{acc}_{k}(h)$. ∎
B.4 Retraining
Figure 6: Theorem 4.3 – Illustration of the proof. Pink line: The expected accuracy of a fixed classifier $h$ is linear in ${\bm{p}}$ (Proposition B.1). Blue line: The overall accuracy of the optimal classifier $h_{\bm{p}}$ is a pointwise maximum over linear functions, and is therefore convex (Proposition B.9). The gradient of $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ over the simplex is the vector of marginal accuracies, showing that the overall accuracy is a potential function for the game (Proposition B.12). Properties of the game’s equilibria then follow from convexity arguments.
Definition B.7 (Retraining prediction game with an optimal predictor).

### v2

@@ D.1 Preliminaries
The set of distribution over a set $A$ is denoted by ${\Delta\left({A}\right)}$, and the set of distributions over $[K]$ is denoted by $\Delta^{K}={\Delta\left({[K]}\right)}={\left\{{{\bm{p}}\mid\sum_{k}p_{k}=1,p_{k}\geq 0\ \ \forall k\in[K]}\right\}}$. The support of a distribution ${\bm{p}}\in\Delta^{K}$ is denoted by $\operatorname*{support}({\bm{p}})={\left\{{k\in[K]\mid p_{k}>0}\right\}}$. We denote the space tangent to the $K$-simplex by $T\Delta^{K}={\left\{{z\in\mathbb{R}^{K}\mid\sum_{k}z_{k}=0}\right\}}$. We denote ${\bm{1}}=\left(1,\dots,1\right)\in\mathbb{R}^{K}$, and denote by ${\bm{\Phi}}={\bm{I}}-{\bm{1}}{\bm{1}}^{T}$ the orthogonal projection matrix from $\mathbb{R}^{K}$ to $T\Delta^{K}$. For $k\in[K]$, we denote by $\mathbf{e}_{k}\in\Delta^{K}$ the singleton categorical distribution supported over $k$, representing a mixture which only contains the distribution $D_{k}$.
@@
We denote the space of features by $\mathcal{X}$, the space of labels by $\mathcal{Y}$, the hypotheses class by $\mathcal{H}\subseteq\mathcal{Y}^{\mathcal{X}}$, and the loss function by $\mathcal{L}:\mathcal{Y}^{2}\to\mathbb{R}$. The expected loss of hypothesis $h\in\mathcal{H}$ under population distribution $D\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ is denoted by $\mathcal{L}_{D}(h)=\mathbb{E}_{(x,y)\sim D}{\left[{\mathcal{L}(h(x),y)}\right]}$. A supervised learning algorithm $\mathcal{A}$ is a possibly-stochastic mapping from a distribution $D\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$ to a classifier $h\sim\mathcal{A}(D)$, formally $\mathcal{A}:{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}\to{\Delta\left({\mathcal{H}}\right)}$.
@@
Each group $k$ is associated with a feature-label distribution $D_{k}\in{\Delta\left({{\mathcal{X}\times\mathcal{Y}}}\right)}$. When the population state is ${\bm{p}}\in\Delta^{K}$, the training state is a mixture distribution denoted by $D_{\bm{p}}=\sum_{k\in[K]}p_{k}D_{k}$. We use a vector subscript notation to distinguish between group distributions (e.g., $D_{k}$) and mixture distributions (e.g., $D_{\bm{p}}$). Given hypothesis $h\in\mathcal{H}$, the expected loss of group $k$ is denoted by $\mathcal{L}_{k}(h)=\mathbb{E}_{(x,y)\sim D_{k}}{\left[{\mathcal{L}(h(x),y)}\right]}$, and the vector of expected group losses is denoted by $\vec{\mathcal{L}}(h)=\left(\mathcal{L}_{1}(h),\dots,\mathcal{L}_{K}(h)\right)$. For ${\bm{p}}\in\Delta^{K}$, the expected loss of $h$ under distribution $D_{\bm{p}}$ is denoted by $\mathcal{L}_{\bm{p}}(h)=\mathbb{E}_{(x,y)\sim D_{\bm{p}}}{\left[{\mathcal{L}(h(x),y)}\right]}$.
@@
We follow the notations of Sandholm [43]. Population games associate each group $k\in[K]$ with a fitness function $F_{k}({\bm{p}})$, which maps a population state ${\bm{p}}\in\Delta^{K}$ to the evolutionary fitness of group $k$. A Nash Equilibrium (NE) of a population game is a population state ${\bm{p}}^{*}$ which satisfies $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k\in[K]}F_{k}({\bm{p}}^{*})$, and every population game admits at least one Nash equilibrium [43, Theorem 2.1.1]. A population game is a potential game if it can be represented as a gradient of some function. Formally, $F:\Delta^{K}\to\mathbb{R}^{K}$ is a potential game if there exists $f:\Delta^{K}\to\mathbb{R}$ such that $F=\nabla f$, where $\nabla f$ is the gradient of $f$ over the simplex (see [43], Section 3.2 for rigorous definitions).
Competition, mutualism, and evolutionary stability.
Coexistence equilibria are fixed points in which two groups or more persist ${\left|{\operatorname*{support}({\bm{p}}^{*})}\right|}\geq 2$. We say that coexistence is competitive if each group’s accuracy at ${\bm{p}}^{*}$ is lower than its accuracy in isolation (i.e., when it is the sole group in the population). Conversely, we say that interactions are mutualistic if every group’s accuracy at ${\bm{p}}^{*}$ exceeds the accuracy it obtains in isolation, indicating that coexistence benefits all groups (see Figure 5 for an illustration). Additionally, we say an equilibrium is evolutionarily stable if it attracts nearby population states (see [43, Section 8.3] for formal definitions).
D.2 Evolutionary Dynamics
Figure 6: Evolutionary dynamics of an oracle classifier in a three-group setting. (Left) Data distributions of groups ${\left\{{A,B,C}\right\}}$. (Right) Phase diagram of replicator dynamics over the probability simplex, induced by the evolutionary prediction game. Background color represents the overall accuracy $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$, vectors represent the flow of replicator dynamics, and markers represent fixed points, with marker color indicating stability. The dynamics have 7 fixed points: One unstable fixed point at the interior of the simplex (unstable coexistence equilibrium), three stable fixed points at the vertices of the simplex (stable equilibria with single-group dominance), and three additional unstable points at the boundaries (unstable restricted equilibria induced by the replicator dynamics).
For our theoretical analysis, we assume that ${\bm{p}}$ evolves in time according to an ordinary differential equation of the form $\dot{{\bm{p}}}=V_{F}({\bm{p}})$, where $V_{F}:\Delta^{K}\to T\Delta^{K}$ is a vector field tangent to the simplex, induced by the fitness function $F({\bm{p}})$, assumed to be continuous (see [43, Theorem 4.4.1]). A common example for such dynamical system is the replicator equation $\dot{p}_{k}=p_{k}(F_{k}({\bm{p}})-\sum_{k^{\prime}}F_{k^{\prime}}({\bm{p}}))$, which corresponds to the tangent vector field $V_{F}({\bm{p}})={\bm{p}}\odot\left(F({\bm{p}})-\sum_{k^{\prime}}p_{k^{\prime}}F_{k^{\prime}}({\bm{p}}){\bm{1}}\right)$, where $\odot$ is the Hadamard product (see Figure 6). The replicator dynamics are a common model of evolutionary dynamics (see Appendix C, and e.g. [49, 26]), and satisfies our core assumptions. Moreover, we also emphasize that our analysis applies to any $V_{F}$ which satisfies the core assumptions below.
Assumptions.
With respect to $V_{F}$, we make the following assumptions:
  • Continuity: We assume that $V_{F}({\bm{p}})$ is continuous with respect to ${\bm{p}}$.
  • Positive correlation: We assume that $V_{F}({\bm{p}})$ is directionally aligned with the fitness vector $F({\bm{p}})$ whenever a population is not at rest, formally $V_{F}({\bm{p}})\cdot F({\bm{p}})>0$ whenever $V_{F}({\bm{p}})\neq{\bm{0}}$.
  • Nash stationarity: We assume that fixed points $V_{F}({\bm{p}}^{*})={\bm{0}}$ correspond to Nash equilibria ${\bm{p}}^{*}$ of the game $F$ (see eq. 3 for the definition of Nash equilibrium).
For a detailed discussion of the positive correlation assumption, see [43, Section 5.2]. We also note that imitative dynamics (such as the replicator equation), may admit additional rest points at restricted equilibria of the population game (formally, states satisfying $F_{k}({\bm{p}}^{*})=F_{k^{\prime}}({\bm{p}}^{*})$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}}^{*})$). However, we note that this does not affect our main results (Theorem 1 and its extensions, and our constructions in Section 5), as any Nash equilibrium is also a restricted equilibrium, and any non-Nash restricted equilibrium corresponds to an unstable fixed point of the imitative dynamics (see [43, Section 8.1]). See Figure 6 for an illustration of Nash and restricted equilibria in a three-group setting.
Time scales.
Rates of convergence can be tuned by scaling the vector field $V_{F}$. For example, replacing $V_{F}$ with $\alpha V_{F}$ for $\alpha>0$ compresses or stretches time by a factor of $\alpha$ while still maintaining our core assumptions. In particular, if ${\bm{p}}^{t}$ satisfies $\dot{\bm{p}}=V_{F}({\bm{p}})$, then ${\bm{q}}^{t}={\bm{p}}^{\alpha t}$ is a solution to $\dot{\bm{q}}=\alpha V_{F}({\bm{q}})$ with the same initial conditions. Thus, our convergence and stability results hold irrespective of whether the dynamics evolve quickly or slowly.
Dynamics induced by potential games.
@@ D.2 Evolutionary Dynamics
Dynamics induced by potential games are closely related to the structure of the potential function $f({\bm{p}})$. We informally state key results that are relevant to our setting: By [43, Theorem 3.1.3], a state ${\bm{p}}^{*}$ is a Nash equilibrium of $F$ if and only if it is an extremum of the potential function $f$. By [43, Lemma 7.1.1] potential function are Lyapunov functions, and by [43, Theorem 7.1.2], dynamics satisfying the positive correlation property converge towards Nash equilibria (restricted Nash equilibria in case of replicator dynamics). Finally, by [43, Theorem 8.2.1], a population state ${\bm{p}}^{*}$ is a stable equilbrium if and only if it is a local maximizer of $f({\bm{p}})$.
D.3 Basic Claims
D.3.1 Linearity
Proposition 1 (Expected loss is linear in ${\bm{p}}$ for a fixed hypothesis).
@@ D.3.1 Linearity
D.3.2 Overall Accuracy Equality
Equilibria of evolutionary prediction games also satisfy a natural fairness criterion. A classifier $h$ satisfies overall accuracy equality if all groups have equal prediction accuracy. In the context of natural selection, we consider this with respect to the groups currently present in the population:
Definition 1 (Overall accuracy equality; [e.g., 51]).
Let $h:\mathcal{X}\to\mathcal{Y}$, and let ${\bm{p}}\in\Delta^{K}$. $h$ satisfies overall prediction equality if $\mathrm{acc}_{k}(h)=\mathrm{acc}_{k^{\prime}}(h)$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}})$.
Under retraining, a different classifier $h\sim\mathcal{A}({\bm{p}})$ is deployed at each time step. It is therefore natural to define this property with respect to the learning algorithm, rather than a single classifier. In the following definition, we take the expectation over the stochasticity of the sampling and learning process to representing the average over time under retraining:
Definition 2 (Overall accuracy equality in expectation).
Let ${\bm{p}}$ be a population state, and let $\mathcal{A}({\bm{p}})$ be a learning algorithm. $\mathcal{A}$ satisfies overall prediction equality in expectation if $\mathbb{E}_{h\sim\mathcal{A}({\bm{p}})}{\left[{\mathrm{acc}_{k}(h)}\right]}=\mathbb{E}_{h\sim\mathcal{A}({\bm{p}})}{\left[{\mathrm{acc}_{k^{\prime}}(h)}\right]}$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}})$.
Proof of Proposition 1.
Let ${\bm{p}}^{*}$ be a Nash equilibrium of the evolutionary prediction game induced by $\mathcal{A}({\bm{p}})$. By eq. (3), it holds that $\operatorname*{support}({\bm{p}}^{*})\subseteq\operatorname*{arg\,max}_{k}F_{k}({\bm{p}}^{*})$, and therefore there exists some $a\in\mathbb{R}$ such that $F_{k}({\bm{p}}^{*})=a$ for all $k\in\operatorname*{support}({\bm{p}}^{*})$. Then by the definition of the evolutionary prediction game (Def. 1), it holds that: $\mathbb{E}_{h\sim\mathcal{A}({\bm{p}})}{\left[{\mathrm{acc}_{k}(h)}\right]}=\mathbb{E}_{h\sim\mathcal{A}(D)}{\left[{\mathrm{acc}_{k^{\prime}}(h)}\right]}=a$ for all $k,k^{\prime}\in\operatorname*{support}({\bm{p}}^{*})$. ∎
D.4 Training Once
We begin with a characterization of evolutionary dynamics for a simple setting in which the classifier is trained only once on the initial distribution, and henceforth kept fixed. This is useful as a first step since by fixing $h=h^{0}$, outcomes can be attributed solely to the induced changes in the population.
We will prove the following theorem:
Theorem 1.
  1. 1.
    Competitive exclusion: ${\bm{p}}^{*}$ is supported on $k^{*}\in\operatorname*{arg\,max}_{k}\mathrm{acc}_{k}(h^{0})$, i.e., initially-fittest groups.
  2. 2.
    Stability: ${\bm{p}}^{*}$ is stable.
  3. 3.
    Accuracy: $\tfrac{\mathrm{d}}{\mathrm{d}t}\mathrm{acc}_{{\bm{p}}^{t}}(h^{0})\geq 0$, i.e., overall accuracy weakly improves over time.
Definition 3 (Train-once prediction game).
@@ D.4 Training Once
Definition 4 (Constant game; [43, Section 3.2.4]).
@@
Fix $h_{0}\in\mathcal{Y}^{\mathcal{X}}$. In the train-once setting, the expected marginal accuracies are constant as a function of ${\bm{p}}$, and therefore $F_{k}$ is a constant game by definition 4. By [43, Proposition 3.2.13], $F_{k}$ a potential game with linear potential $f({\bm{p}})=-\vec{\mathcal{L}}(h^{0})\cdot{\bm{p}}$. By [43, Theorem 3.1.3], the set of Nash equilibria is the set of local maximizers of $f({\bm{p}})$, which is a convex combination over the set $\operatorname*{arg\,min}_{k\in[K]}\mathcal{L}_{k}(h^{0})$ due to linearity. Hence:
@@
At equilibrium ${\bm{p}}^{*}$, it holds that $\mathcal{L}_{{\bm{p}}^{*}}=\min_{k\in[K]}\mathcal{L}_{k}(h^{0})$. Apply proposition 1 to obtain:
@@
Finally, observe that each ${\bm{p}}^{*}$ is stable as a maximizer of the linear potential function [43, Theorem 8.2.1]. ∎
Proof of Theorem 1.
Theorem 1 is a special case of Proposition 2 for the loss function: $\mathcal{L}_{k}(h)=-\mathrm{acc}_{k}(h)$. ∎
D.5 Retraining
Figure 7: Illustration of the proof of Theorem 1. Pink line: The expected accuracy of a fixed classifier $h$ is linear in ${\bm{p}}$ (Proposition 1). Blue line: The overall accuracy of the oracle classifier $h_{\bm{p}}$ is a pointwise maximum over linear functions, and is therefore convex (Proposition 3). The gradient of $\mathrm{acc}_{\bm{p}}(h_{\bm{p}})$ over the simplex is the vector of marginal accuracies, showing that the overall accuracy is a potential function for the game (Proposition 4). Properties of the game’s equilibria then follow from convexity arguments.
Definition 5 (Retraining prediction game with an oracle predictor).